OpenStudy (anonymous):

multiply, then simplify where possible. (square root of 5)^2

OpenStudy (anonymous):

is the problem\[\sqrt{5}^{ 2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if the squared is NOT inside the radical, the ans is 5

OpenStudy (anonymous):

and if the square is inside the radical it is 5 as well.

OpenStudy (anonymous):

\[\sqrt{5^2}=\sqrt{25}=5\]